Starset, Inc., has a target debt-equity ratio of 0.77. Its WACC is 11 percent, and the tax rate is 32 percent.

a. If the company's cost of equity is 16.5 percent, what is the pretax cost of debt?

b. If instead you know that the aftertax cost of debt is 6.6 percent, what is the cost of equity?
Answer #1

Debt-equity ratio=debt/equity

Hence debt=0.77*equity

Let equity be $x

Debt=$0.77x

Total=$1.77x

WACC=Respective cost*Respective weight

a.

11=(x/1.77x*16.5)+(0.77x/1.77x*Cost of debt)

11=9.3220339+(0.77x/1.77x*Cost of debt)

Cost of debt=(11-9.3220339)*1.77/0.77

=3.85714285%(Approx)

Pretax cost of debt=Cost of debt/(1-tax rate)

=3.85714285/(1-0.32)

=5.67%(Approx).

b.

11=(x/1.77x*Cost of equity)+(0.77x/1.77x*6.6)

11=(x/1.77x*Cost of equity)+2.87118644

Cost of equity=(11-2.87118644)*1.77

=14.388%